OpenStudy (anonymous):

Need help with Absolute Value Equations for Algebra 1 Fan And Medal

OpenStudy (jdoe0001):

posting them will help

OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21/4 , x = −11/4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (jdoe0001):

\(\bf -4|x + 5| = -16\implies |x + 5| =\cfrac{-16}{-4}\implies |x + 5| =4\\ \quad \\ |x + 5| =4\implies \begin{cases} +(x + 5) =4\\ \quad \\ \bf -(x + 5) =4 \end{cases}\) so you'd get 2 values for "x", just solve each scenario

OpenStudy (anonymous):

I keep on getting -5 for both of them

OpenStudy (jdoe0001):

\(\bf |x + 5| =4\implies \begin{cases} +(x + 5) =4\implies x+5=4\\ \quad \\ \bf -(x + 5) =4\implies -x-5=4 \end{cases} \)

OpenStudy (jdoe0001):

neither one is -5 btw

OpenStudy (anonymous):

I got -1 and 9 , Is that right ?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

well.. -1 and ... -9

OpenStudy (anonymous):

Can you help me with this one too please ? I just want to make sure that I understand it . Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (jdoe0001):

well. you do the same thing, you isolate the absolute value expression, and then solve the 2 scenarios \(\bf |x + 2| + 16 = 14\implies|x + 2| =14-16\implies|x + 2|=-2\\ \quad \\ |x + 2|=-2\implies \begin{cases} +(x + 2)=-2\\ \quad \\ \bf -(x + 2)=-2 \end{cases}\)

OpenStudy (anonymous):

-4 and 0 ?

OpenStudy (anonymous):

Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (jdoe0001):

well... based on the scenarios of the previous ones, what do you think? how many solutions do each one have?

OpenStudy (anonymous):

2 ?

OpenStudy (jdoe0001):

:)